# Thread: A Logic-Based Data Organization Paradigm for How-To Question Answering with Retrieval Augmented Generation

### v1

@@ Abstract
Current question answering systems leveraging retrieval augmented generation perform well in answering factoid questions but face challenges with non-factoid questions, particularly how-to queries requiring detailed step-by-step instructions and explanations. In this paper, we introduce Thread, a novel data organization paradigm that transforms documents into logic units based on their inter-connectivity. Extensive experiments across open-domain and industrial scenarios demonstrate that Thread outperforms existing data organization paradigms in RAG-based QA systems, significantly improving the handling of how-to questions.
@@ 1 Introduction
Dating back to ancient philosophy, Aristotle’s Nicomachean Ethics Crisp (2014) introduced the concept of “5Ws and 1H” questions to comprehensively understand actions and the circumstances surrounding them, where the “5Ws” represent What, Why, When, Where, and Who, and the “1H” stands for How. These questions serve as foundational tools for addressing various aspects of events and human behavior, spanning from philosophical discourse to daily life questions. In our everyday experiences, the 5Ws and 1H questions cover a wide range of questions across different contexts Kipling (2018); Wikipedia (2024). While the 5Ws questions are typically factoid questions used to gather factual information about an event or situation111“Why” question can sometimes delve into explanations and motivations, thus blurring the line between factoid and non-factoid questions., the “How” question (or referred to in this paper as “How-To” questions) often delves into processes, methods, or explanations, which may involve more interpretation or analysis rather than straightforward facts, tending to be non-factoid questions. Current research has achieved remarkable success in factoid question answering (factoid QA) Nguyen et al. (2016); Rajpurkar et al. (2018); Jiang et al. (2019); Kwiatkowski et al. (2019); Stelmakh et al. (2022). However, research on non-factoid QAs Bolotova et al. (2022); Rogers et al. (2023), particularly those involving “How-To” questions, remains relatively underexplored. Existing QA systems often struggle to provide detailed step-by-step answers to how-to questions akin to human comprehension, highlighting a significant gap in current AI capabilities Krishna et al. (2021); Soleimani et al. (2021).
Figure 1: Two types of How-To questions.
In order to better address how-to questions, we explicitly propose the definitions of two types of how-to questions: Linear How-To Questions and Dynamic How-To Questions. Linear how-to questions involve a fixed sequence of steps that do not require feedback or decision points based on intermediate outcomes. These steps are predefined and must be followed in a specific order to achieve the desired outcome, making them suitable for straightforward, procedural tasks Merrill (2012). For example, a question like “How to cook spaghetti?” would be answered with a step-by-step procedure as shown in Figure 1. On the other hand, dynamic how-to questions require a more complex approach where each step may lead to different actions based on the results of previous steps. This type involves conditional decision points and may require iteration, making the process adaptive and responsive to varying circumstances Newell et al. (1972); Wang and Ruhe (2007). As shown in Figure 1, initial steps are followed by different paths, where each path requiring different steps and potentially iterative troubleshooting.
The nature of how-to QA inherently follows the decision-making processes central to problem-solving Polya (2004) and human learning in cognitive science Learn (2000). This requires a structured approach to break down and solve tasks step by step. However, such a natural extension of human decision-making makes how-to questions pose significant challenges for QA systems leveraging Retrieval Augmented Generation (RAG). In the era of Large Language Models (LLMs), RAG represents a powerful technique to enhance the capability of LLMs by incorporating external knowledge. Current RAG systems mainly concentrate on improving the retrieval quality or adaptability (Shao et al., 2023; Trivedi et al., 2023; Jiang et al., 2023; Asai et al., 2023). Despite their success on factoid questions, they still struggle to handle how-to questions, particularly dynamic how-to questions. The root cause of this problem lies in the data organization they use during the retrieval stage. The common practice of splitting documentation into chunks and indexing these chunks for retrieval can disrupt the inherent connections within the text. Although recent works have discussed the granularity of chunks Chen et al. (2023); Gao et al. (2023), they still adhere to the chunk-based222“Chunk” refers to a general document splitting paradigm including chunks, sentences, phrases, etc. data organization paradigm. This approach can fragment the solution to a how-to question into several chunks, leading to scenarios where the retriever fails to gather all relevant chunks based on semantic or lexical similarity. Consequently, the answers may lack coherence and fail to maintain the logical sequence required for accurately addressing how-to questions.
To facilitate current RAG systems on how-to questions, in this paper, we propose a new logic-based data organization paradigm called Thread, which is compatible with RAG systems. Thread explores the inner connections within the documents, splitting the documents into more structured and loosely interconnected logic units (LUs). Given a how-to question, the retriever gets the relevant LUs with its indexed header, and the linker of retrieved LUs will automatically link to other LUs, till obtaining enough information to answer the how-to question. It makes the process of answering how-to questions like “Pulling on the thread, the whole mystery started to unravel like a sweater.” Garcia and Stohl (2011)
@@ 2 Related Work
2.1 Data Organization Paradigm in RAG
The data organization process is a critical pre-stage of RAG methods where documents are processed and segmented following certain data organization paradigms. The most common data organization paradigm is splitting documents into retrieval units Gao et al. (2023). These retrieval units vary in granularity such as phrases, sentences, propositions Chen et al. (2023), chunks Kamradt (2024), etc. Coarser-grained units contain more information but introduce redundant noise, while finer-grained units have lower semantic integrity and often require retrieving more units to gather comprehensive information. However, the chunk-based data organization paradigm ignores the logical and relational connections between chunks, potentially disrupting the inherent logic flow in documents. Another paradigm constructs documents into knowledge graphs (KG), where retrieval units include entities, triplets, etc. Gaur et al. (2022); Sen et al. (2023); He et al. (2024); Wang et al. (2024); Edge et al. (2024). However, these approaches emphasize semantic/lexical similarities between retrieval units, their success in factoid QA is limited when applied to how-to QA. This limitation arises because how-to QA demands logical connections between retrieval units that extend beyond mere semantic or lexical similarities.
2.2 Information Retrieved by RAG
The effectiveness of RAG methods depends on the generator’s ability to utilize retrieved information and the quality and quantity of that information. Insufficient question-relevant information can cause hallucination in LLM-based generators Li et al. (2023); Zhang et al. (2023), making it crucial to improve the retrieval process. Traditional one-round retrieval methods Guu et al. (2020); Lewis et al. (2020) often fail to gather all necessary information due to their reliance on the similarity between query and retrieval units Gan et al. (2024). Advanced RAG methods use query rewriting and expansion Peng et al. (2024); Shao et al. (2023); Trivedi et al. (2023); Kim et al. (2023) or iterative retrieval Shao et al. (2023); Jiang et al. (2023); Asai et al. (2023) to collect more information. However, these approaches still struggle with dynamic how-to questions, which require making next-step decisions based on the current retrieved units, unless the current retrieved units contain clues that lead to the next step. The main issue is the lack of connections between retrieval units, which prevents effective retrieval and the gathering of sufficient information for answering how-to questions.
@@ 3.1 Logic Unit: Retrieval Unit of Thread
We propose a new data organization paradigm called Thread. It is a knowledge base composed of discrete but interconnected retrieval units called Logic Units (LUs). Each consists of the following:
Prerequisite. The prerequisite component acts as an information supplement, providing the necessary context to understand the LU. For example, an LU may include domain-specific terminology such as entities or abbreviations. The prerequisite explains these terms and can generate new queries to retrieve LUs with more detailed information. Without this context, passing these LUs to an LLM-based generator could lead to hallucinations. Additionally, the prerequisite can function as an LU filter, containing constraints that must be met before the LU is considered in answer generation. This filtering ensures only relevant LUs are retrieved. Tags are a specific example of such prerequisites, quickly filtering out irrelevant LUs in the retrieval stage.
Header. The header summarizes the LU or describes the intention it aims to address, depending on the type of LU (refer to §3.2). For example, the header could be the name of a terminology if LU describes a terminology; if the LU describes actions to resolve a problem, the header describes the intent or the problem LU aims to resolve. The header is used for indexing, similar to traditional RAG processes, and serves as the key for retrieving the LU based on a query.
@@
Linker. The linker acts as a bridge between logic units. It is used to generate new queries for the next-round retrieval. In traditional iterative RAG methods, query generation relies on previous retrieval units, but these units often lack direct clues to retrieve new information. In Thread, the linker provides the necessary information for generating these queries. For example, in dynamic how-to questions, the linker specifies multiple possibilities after taking the action in the LU body, guiding the retrieval of the next-step LU. The format of the linker varies depending on the LU type; it can be a query connecting to retrieve other LUs or an entity relationship. The edge of knowledge graph in traditional factoid QA is a special linker enabling navigation between related entities. If no further LUs are connected, the linker remains empty, isolating the current LU.
@@ 3.2 Logic Unit Type
When converting documents into logic units (LUs), there are multiple types of LUs. Below are the common LU types identified in our experiments [^3] These LU types are summarized from our practice in experiments with industrial and public datasets. There may be additional LU types depending on the specific scenario.:
@@
Terminology. This type provides detailed explanations of domain-specific terminology. For example, terms may share the same name or abbreviation in the LU header but have different meanings in the LU body. The prerequisite in terminology LUs includes scenarios where the terminology typically appears. Terminology LUs generally have empty linkers unless they refer to extended terminology that depends on it.
@@ 3.3 Thread: LU-based Knowledge Base
Figure 2: The process of Thread construction.
In practice, documentation is often unstructured and varies in format and style. Our approach to converting documentation into Thread involves a two-stage process to obtain LUs. We believe that with advancements in LLMs, this process could be streamlined into a single stage, enabling high-fidelity LU conversion with minimal hallucination.
Documentation Reformulation (Optional). This stage is optional, depending on the quality of the documentation. For example, in software engineering, Troubleshooting Guides (TSGs) often lack readability and detail, negatively impacting productivity and service health Shetty et al. (2022). Due to varying document styles, where some are clearly outlined and others are disordered, we avoid directly extracting LUs from the original documents. Instead, we first reformulate these documents into structured formats. Leveraging LLMs for this task, we enhance the LLMs’ in-domain understanding by providing search capabilities and domain-specific context. This is followed by a refinement step to prevent overlooking details or hallucinating information. Figure 2(a) shows the reformulation stage. The reformulation stage is unnecessary for well-written documents like product help docs, which typically follow a linear how-to format. Table C in the Appendix C shows an example of this process with prompts.
@@
As we extract and merge LUs, the collection of LUs from all documents forms the knowledge base, i.e., Thread. This LU-based knowledge base serves as an essential component compatible with the current RAG-based QA system and even makes it possible for automation.
Figure 3: The QA system integrated with Thread. It retrieves relevant LUs based on query-Header similarity and filters out LUs that do not meet the current prerequisites. The selected LUs are passed to LLMs to generate actions for execution. After execution, the Linker matches results and generates a new query for the next retrieval iteration.
@@ 3.4 Integrate Thread with QA System
To demonstrate how Thread works, we use dynamic how-to questions as an example. Figure 3 shows the QA system incorporating our Thread data organization paradigm. The Retriever and LLM-based Generator are components from the original RAG-based QA system. LUs are indexed by their Headers. When an initial how-to question is submitted, the Retriever identifies the top-K most relevant LUs based on query-Header similarity. The Selector then checks the prerequisites of these LUs and filters out those that do not meet the current prerequisites, derived from the initial question or any available chat history. If no current prerequisite is provided, the system can prompt the user, e.g., “Before doing …, have you tried …?”, to obtain the current prerequisite for LU filtering. After selection, the Body of the LUs is fed into the LLM-based generator to produce an answer. If an execution engine is available, actions can be executed automatically; otherwise, the answer/action is presented to the user for manual execution. Once the action is executed, the Linker matches one of the possible outcomes and generates a new query for the next retrieval round.
Unlike traditional QA systems, the QA system with Thread can potentially be fully or semi-automated if an execution engine is present. Compared to manually designed automation pipelines, Thread enables greater automation and flexibility where updating LUs will automatically update the semi-automated system (see §3.3).

### v2

@@ Abstract
Recent advances in retrieval-augmented generation have significantly improved the performance of question-answering systems, particularly on factoid ‘5Ws’ questions. However, these systems still face substantial challenges when addressing ‘1H’ questions, specifically how-to questions, which are integral to decision-making processes and require dynamic, step-by-step answers. The key limitation lies in the prevalent data organization paradigm, chunk, which divides documents into fixed-size segments, and disrupts the logical coherence and connections within the context. To overcome this, in this paper, we propose Thread, a novel data organization paradigm aimed at enabling current systems to handle how-to questions more effectively. Specifically, we introduce a new knowledge granularity, termed ‘logic unit’, where documents are transformed into more structured and loosely interconnected logic units with large language models. Extensive experiments conducted across both open-domain and industrial settings demonstrate that Thread outperforms existing paradigms significantly, improving the success rate of handling how-to questions by 21% to 33%. Moreover, Thread exhibits high adaptability in processing various document formats, drastically reducing the candidate quantity in the knowledge base and minimizing the required information to one-fourth compared with chunk, optimizing both efficiency and effectiveness.
@@ 1 Introduction
Question answering (QA) is a foundational research topic in human-machine interaction (Allam & Haggag, 2012). Among the most advanced techniques, Retrieval-augmented generation (RAG) organizes external documents into fixed-size chunks and retrieves relevant knowledge to enhance QA systems (Shao et al., 2023; Trivedi et al., 2023; Jiang et al., 2023; Asai et al., 2023). This routine is particularly effective in handling the ‘5Ws’ questions, which are typically factoid requiring factual information about specific details (Yang et al., 2018; Jiang et al., 2019; Kwiatkowski et al., 2019; Stelmakh et al., 2022), such as ‘When is Shakespeare’s birthday?’. Specifically, the data organization paradigm only needs to provide chunks containing the relevant knowledge, e.g. triples or documents about the topic entity, to the RAG systems to facilitate the ‘5Ws’ questions. However, the ‘1H’ questions, derived from Aristotle’s Nicomachean Ethics (in this paper as ‘how-to’ questions, Crisp 2014), remain largely underexplored [^1] The ‘5Ws’ represent What, Why, When, Where, and Who, and the ‘1H’ stands for How. The ‘Why’ question can sometimes delve into explanations and motivations.. These questions are in high demand in practical applications such as teaching us how to write code to achieve specific goals.
Central to problem-solving (Polya & Pólya, 2014) and human learning in cognitive science (Learn, 2000), how-to questions are inherently more complex, often involve processes that require interpretation and analysis (Deng et al., 2023b). For example, correctly answering the question shown in Figure 1How to diagnose and fix a performance issue in a web application?’ involves a multi-step decision-making process, i.e., firstly checking the server load and response time, then optimizing server configuration or indexes and queries with user feedback. Such a dynamic nature necessitates the RAG systems incrementally guide users through each step, adapting to specific circumstances and providing precise and logical information. However, the prevalent data organization paradigm, chunk222‘Chunk’ refers to a general document splitting paradigm including chunks, sentences, phrases, etc. (Splitter, 2023; Chen et al., 2023; Gao et al., 2023) leaves current RAG systems (Asai et al., 2023; Shao et al., 2023) struggling to solve such questions. This chunk-based paradigm disrupts the logical flow and coherence of the content due to its inability to effectively represent the internal logic. Consequently, it returns excessive information during inference, making it difficult to maintain continuity between sequential steps, leading to fragmented responses for how-to questions and posing significant challenges for users. Overcoming these limitations necessitates a paradigm shift that accounts for the logical structure and stepwise nature of how-to questions, enabling more precise and coherent information retrieval tailored to specific needs.
Figure 1: An example of how-to questions with its decision-making process. We omit details such as the actions to check the server load and response time due to limited space.
In this paper, we propose Thread, a new logic-based data organization paradigm designed to enhance the handling of how-to questions. The name Thread evokes the idea of ‘Pulling on the thread, the whole mystery started to unravel like a sweater.’ (Garcia & Stohl, 2011). Specifically, we introduce a new knowledge granularity named ‘logic unit’, comprising five key components and four different types (see §3.1 and §3.2). We employ a two-stage process depicted in §3.3 to extract logic units (LUs) from documents. The first, optional stage is reformulating the original documents depending on their format and style, and the second focuses on extracting and merging logic units. In this way, Thread captures connections within the documents, breaking them into more structured and loosely interconnected logic units. When answering how-to questions, the system integrated with Thread enables a dynamic interaction manner. First, it retrieves relevant LUs based on their indexed headers. Then, the body of the selected LU provides the necessary content to generate responses for the current step. With user feedback, the linker in LU dynamically connects to other LUs, allowing the system to adapt its responses until the how-to question is comprehensively addressed.
We evaluate the effectiveness of Thread by conducting experiments in two open-domain, web navigation (Deng et al., 2023a) and Wikipedia Instructions (Koupaee & Wang, 2018), and one industrial setting, Incident Mitigation (Shetty et al., 2022). Experimental results demonstrate that while existing paradigms struggle with how-to questions, Thread excels at handling consecutive steps and consistently outperforms them, especially in real-world incident mitigation scenarios, with an improvement of Success Rate ranging from 21.05% to 33.33%. In addition, Thread demonstrates great efficiency and effectiveness in processing various document formats, significantly reducing the number of retrieval units to 20% and the length of tokens needed for generation to 22.60%. The main contributions of this paper include:
  • We highlight the challenges faced by current RAG systems in addressing how-to questions. To address the limitation of the chunk-based paradigm, we propose Thread, a novel data organization paradigm that transforms original documents into structured, interconnected logic units.
  • Integrated with Thread, our system follows a dynamic interaction mode, guiding users incrementally through each step and adapting to their specific circumstances. Our system also brings more possibilities for an automation pipeline, solving how-to questions more efficiently.
  • Experimental results show that Thread significantly outperforms existing data organization paradigms across three scenarios. Additionally, Thread requires less information, while effectively handling various document formats and substantially reducing the knowledge base size.
@@ 2 Related Work
Data Organization Paradigm in RAG
The data organization process is a critical pre-stage of RAG methods where documents are processed and segmented following certain data organization paradigms. The most common data organization paradigm is splitting documents into retrieval units (Gao et al., 2023). These retrieval units vary in granularity such as phrases, sentences, propositions (Chen et al., 2023), chunks (Kamradt, 2024), etc. Coarser-grained units contain more information but introduce redundant noise, while finer-grained units have lower semantic integrity and often require retrieving more units to gather comprehensive information. However, the chunk-based data organization paradigm ignores the logical and relational connections between chunks, potentially disrupting the inherent logic flow in documents. Another paradigm constructs documents into knowledge graphs (KG), where retrieval units include entities, triplets, etc. (Gaur et al., 2022; Sen et al., 2023; He et al., 2024; Wang et al., 2024; Edge et al., 2024). These approaches emphasize semantic/lexical similarities between retrieval units, their success in factoid questions is limited when applied to how-to questions. This limitation arises because how-to questions demand logical connections between retrieval units that extend beyond mere semantic or lexical similarities.
Information Retrieved by RAG
The effectiveness of RAG methods depends on the generator’s ability to utilize retrieved information and the quality and quantity of that information. Insufficient question-relevant information can cause hallucinations in LLM-based generators (Li et al., 2023; Zhang et al., 2023), making it crucial to improve the retrieval process. Traditional one-round retrieval methods (Guu et al., 2020; Lewis et al., 2020) often fail to gather all necessary information due to their reliance on the similarity between query and retrieval units (Gan et al., 2024). Advanced RAG methods use query rewriting and expansion (Shao et al., 2023; Trivedi et al., 2023; Kim et al., 2023) or iterative retrieval (Shao et al., 2023; Jiang et al., 2023; Asai et al., 2023) to collect more information. However, these approaches still struggle with how-to questions, which require making next-step decisions based on the current retrieved units, unless the current retrieved units contain clues that lead to the next step. The main issue is the lack of connections between retrieval units, which prevents effective retrieval and the gathering of sufficient information.
@@ 3 Methodology
In this section, we introduce ‘logic unit’ with its components and types, then explain the process of constructing the Thread knowledge base, followed by an illustration of how Thread integrates with existing systems.
@@ 3.1 Logic Unit: Retrieval Unit of Thread
We propose a new knowledge granularity called ‘logic unit’. Different from the chunk-based paradigm, LU comprises specially designed components, especially for bridging consecutive steps when addressing how-to questions, considering the internal logic and coherence of documents.
Prerequisite. The prerequisite component acts as an information supplement, providing the necessary context to understand the LU. For example, an LU may include domain-specific terminology such as entities or abbreviations. The prerequisite explains these terms and can generate new queries to retrieve LUs with more detailed information. Without this context, passing these LUs to an LLM-based generator could lead to hallucinations. Additionally, the prerequisite can function as an LU filter, containing constraints that must be met before the LU is considered in answer generation. This filtering ensures only relevant LUs are retrieved. As shown in step LU in Figure 2, accessing the server monitor is a prerequisite for checking the server load.
Header. The header summarizes the LU or describes the intention it aims to address, depending on the type of LU (refer to §3.2). For example, the header could be the name of a terminology if LU describes a terminology; if the LU describes actions to resolve a problem, the header describes the intent or the problem LU aims to resolve. Different from chunk that indexes the entire content, we use the header for indexing which serves as the key for retrieving the LU based on a query.
@@
Linker. The linker acts as a bridge between logic units, enabling the dynamic process of how-to questions. Unlike the chunk-based paradigm, which relies on previous retrieval units that often lack direct clues, the linker in Thread provides necessary information to generate new queries for subsequent retrieval. In Figure 2, the linker of ‘checking server’ specifies multiple possibilities after taking the action in LU body, guiding the retrieval of the next-step LU. Its format varies by LU type, serving as either a query for retrieving other LUs or an entity relationship. The edge of knowledge graph in traditional factoid questions is a special linker enabling navigation between related entities. When no further LUs are connected, the linker remains empty, isolating the current LU.
@@
Figure 2: The construction process of Thread. The left part depicts the two-stage process including reformulating documents into structured ones and extracting and merging logic units, the right part instantiates the extraction by providing an example.
@@ 3.2 Logic Unit Type
When converting documents into logic units, we distinguish how-to questions into ‘linear’ and ‘dynamic’ types, where linear how-to questions often involve a fixed sequence of steps that do not require feedback or decision points based on intermediate outcomes. Below are the common LU types identified in our experiments [^3] These LU types are summarized from our practice in experiments with industrial and public datasets. There may be additional LU types depending on the specific scenario., showing Thread is versatile enough to handle diverse types:
@@
Terminology. This type provides detailed explanations of domain-specific terminology. For example, terms may share the same name or abbreviation in the LU header but convey different meanings in the LU body. The prerequisites for terminology LUs describe scenarios where the terminology typically appears and linkers are usually empty unless referencing extended terminology that depends on it.
@@ 3.3 Thread: LU-based Knowledge Base
In practice, documentation is often unstructured and varies in format and style. Our approach to converting documentation into Thread involves a two-stage process to obtain LUs, as shown in the left part of Figure 2. We believe that with advancements in LLMs, this process could be streamlined into a single stage, enabling high-fidelity LU conversion with minimal hallucination.
Documentation Reformulation (Optional). This stage is optional, depending on the quality of the documentation. For example, in software engineering, Troubleshooting Guides (TSGs) often vary in format, include diverse types of information, and lack readability and detail, negatively impacting productivity and service health (Shetty et al., 2022). Due to such format, where some are clearly outlined and others are disordered, we avoid directly extracting LUs from original documents. Instead, we first reformulate these documents into structured formats. Leveraging LLMs for this task, we enhance the LLMs’ in-domain understanding by providing search capabilities and domain-specific context. This is followed by a refinement step to prevent overlooking details or hallucinating information. Figure 2 (a) shows the reformulation stage and the right represents with symbols. But it is unnecessary for well-written documents like product help docs, which typically follow a linear how-to format. We illustrate the process in Table LABEL:tab:example_lu by example and provide prompts in Appendix B.
@@
As we extract and merge LUs, the collection of LUs from all documents forms the knowledge base, Thread. This LU-based knowledge base serves as an essential component compatible with the current RAG system and even makes it possible for automation.
@@ 3.4 Integrate Thread with QA System
To demonstrate how Thread works, we use dynamic how-to questions as an example. Figure 3 shows the RAG system incorporating our Thread data organization paradigm. The Retriever and LLM-based Generator are inherited from the original RAG system. LUs are indexed by their Headers. When an initial how-to question is submitted, the Retriever identifies the top-K most relevant LUs based on query-header similarity. The Selector then checks the prerequisites of these LUs and filters out those that do not meet the current prerequisites, derived from the initial question or any available chat history. If no current prerequisite is provided, the system can prompt the user, e.g., ‘Before doing …, have you tried …?’, to obtain the current prerequisite for LU filtering. After selection, the body of the LUs is fed into the LLM-based generator to produce an answer. If an execution engine is available, actions can be executed automatically; otherwise, the answer/action is presented to the user for manual execution. Once the action is executed, the Linker matches one of the possible outcomes and generates a new query for the next retrieval round.
Unlike traditional RAG systems, Thread-enabled systems can be potentially fully or semi-automated when integrated with execution engines. This integration offers greater automation and flexibility, as updating LUs automatically updates the system, compared to manually designed pipelines (see §3.3).
Figure 3: The RAG system integrated with Thread. It retrieves relevant LUs based on query-Header similarity and filters out LUs that do not meet the current prerequisites. The selected LUs are passed to LLMs to generate actions for execution. After execution, the Linker matches results and generates a new query for the next retrieval iteration.